# On the Theory of Bulk Viscosity of Cold Plasmas

### v2

@@ Abstract
Solving the kinetic equation for ionization-recombination processes in cold plasmas for temperatures much lower than the first ionization potentials, we derive an explicit expression for the bulk viscosity. We obtain that bulk viscosity can be many order of magnitude bigger than the shear viscosity. Our result for the relaxation time reveals that the Mandelstam-Leontovich approximation for the frequency dependence of the bulk viscosity is in practice an exact result for the cold plasmas. The illustrative numerical examples correspond to the plasma cocktail of the solar chromosphere at the height of the minimal polytropic index. The possible application for the acoustic heating of the inner solar atmosphere up to the transition region is shortly discussed together with the evaluation to confirm the theory by laboratory plasmas.
@@ II.2 The Saha equation
Often the corresponding concentration is used [13, Sec. 104] and it is straightforward to change the notations
@@
Within the framework of the up to now introduced notations, the Saha equation for the equilibrium concentrations reads in its standard form [13, Eq. (104.2)]
@@ II.3 Thermodynamic variables
Knowing these thermodynamic potentials, we can easily calculate heat capacity per unit mass from the derivatives [13, Sec. 16]
$\mathcal{C}_{p}$ $\equiv\left(\dfrac{\partial w}{\partial T}\right)_{\!\!\rho},$ (19)
@@
The sound velocity at evanescent frequency can be calculated by the standard thermodynamic derivatives [11, Eq. (A9)]
@@
In the formulae above, the overline denoting thermal equilibrium values have been omitted again. The velocity $c_{0}$ is the sound velocity at low frequency for which during the adiabatic compression of the plasma, the ionization degrees can follow the equilibrium Saha equation. The general formula above has been solved for pure hydrogen plasma [14, Eq. 96] and [15, Eq. 17]
where $\iota\equiv I/T$, $I=13.6$ eV is the hydrogen ionization potential and $c_{p}\equiv 5/2$ and $c_{v}\equiv 3/2$ are the mono-atomic heat capacities at constant pressure and volume, and here $\alpha\equiv\overline{\alpha}$ is the equilibrium ionization degree
$\alpha\equiv\frac{n_{e}}{n_{p}+n_{0}}=\frac{n_{e}}{n_{\rho}}\in(0,1)\,$, (25)
$n_{e}=n_{p}=\alpha n_{\rho},\quad n_{0}=(1-\alpha)\,n_{\rho},\quad$ (26)
$n_{\mathrm{tot}}=n_{e}+n_{p}+n_{0}=(1+\alpha)\,n_{\rho},$ (27)
where $n_{e}$, $n_{p}$ and $n_{0}$ are the concentrations of electrons, protons and neutral atoms respectively, and we denote the sum of the concentrations of the protons and the neutrals by $n_{\rho}=n_{0}+n_{p}$. This formula can be generalized for arbitrary low temperature cocktail of alkaline metal and noble gas [11, Eq. (B1)], (for example Na-Ne plasma)
$\gamma_{0}=\frac{c_{p}[2+\overline{a}_{\mathrm{nbl}}(2-\alpha)]+(1-\alpha)\alpha(c_{p}+\iota)^{2}}{c_{v}[2+\overline{a}_{\mathrm{nbl}}(2\!-\!\alpha)]+(1\!-\!\alpha)\alpha\left[(c_{v}+\iota)^{2}+c_{v}\right]},$ (28)
where
@@ III Wannier ridge resonances and Wannier near threshold ionization cross-section
It’s remarkable that the exact solution for the scattering in the Coulomb potential coincides with both Born and classical cross-section approximations. The physics is however, absolutely different. At high energies we have slightly perturbed plane waves while the classical approximation is applicable when the velocity $v$ of the scattering electron is much smaller than the Bohr velocity
@@
The numerical values we use $\mathrm{w}\approx 1.18\sim 1$ and $C_{\mathrm{W}}\approx 2.7$ are obtained from the experimental study [20, Fig. 6].
Performing [14, Eq. (5)] the integration over equilibrium Maxwell distribution
@@ IV.4 Generalized polytropic index for cold plasmas
As test for the working programs we have to check that for very low and very high frequencies the polytropic index is real only. For high frequencies it reaches the atomic value $\gamma_{a}$ corresponding to negligible chemical reactions. For low frequencies we have a sophisticated thermodynamic expression Eq. (23) [11, Eq. (8)]
@@ V.1 Mandelstam-Leontovich and Cole-Cole approximation
see also [14, Eq. 99].
The excellent accuracy of the ML fit with fitting parameter time-constant $\tau$ gives a simple interpretation of the both low- and high-frequency limits; we have a simple Padé approximation. We can consider the oscillation of the density as a perturbation and the pressure as a linear response. According to the causality principle, the generalized susceptibility $\hat{p}/\hat{\rho}\propto\hat{\gamma}(\omega)$ must be analytical in the upper semi-plane of the complex frequency $\Im(\omega)>0$ and just along the imaginary semi-axis $\omega=\mathrm{i}\tilde{s}$ with $\tilde{s}>0$ the compressibility must be real and positive. In the lower imaginary semi-plane $\hat{\gamma}_{\mathrm{rel}}^{\mathrm{{ML}}}(\omega)$ has a zero at $\omega=-\mathrm{i}/\tau$ and a pole at $\omega=-\mathrm{i}\Gamma_{0,\infty}/\tau$ in agreement with the general causality principle [13, Sec. 123 Generalized susceptibility].
@@
$A=\int_{0}^{h_{\mathrm{TR}}}2k^{\prime\prime}(x)\mathrm{d}x$ (122)
@@
$\zeta_{0}=-\overline{p}\lim_{\omega\rightarrow 0}\frac{\gamma^{\prime\prime}(\omega)}{\omega}{\color[rgb]{0,0,1}=-\overline{p}\left.\frac{\mathrm{d}\gamma^{\prime\prime}}{\mathrm{d}\omega}\right|_{\omega=0}.}$ (131)
@@
VI Hydrogen-helium cocktail analytical solution
@@
$\overline{\mathcal{N}}_{\mathrm{tot}}=1+\overline{\alpha}+\overline{a}_{\mathrm{He}},\qquad\overline{c}_{e}=\frac{\overline{\alpha}}{1+\overline{\alpha}+\overline{a}_{\mathrm{He}}},$ (149)
@@
The solution of Saha equation for the hydrogen degree of ionization $\overline{\alpha}(T,\overline{n}_{\rho})$ is given in Ref. [14]. Actually, this is an applicable approximation for the solar chromosphere but the derived formulas are applicable for any alkali-noble cocktail, Na-Ne plasma or arbitrary cold pure element plasma.

### v3

# arXiv:2511.14790v3[physics.plasm-ph] 02 Mar 2026

@@ Abstract
Solving the kinetic equation for ionization-recombination processes in cold plasmas for temperatures much lower than the first ionization potentials, we derive an explicit expression for the bulk viscosity. We obtain that bulk viscosity can be many order of magnitude bigger than the shear viscosity. Our result for the relaxation time reveals that the Mandelstam-Leontovich approximation for the frequency dependence of the bulk viscosity is in practice an exact result for the cold plasmas. The illustrative numerical examples correspond to the plasma cocktail of the solar chromosphere at the height of the minimal polytropic index. We systemize also explicit formulae for the thermodynamics of the binary alkali-noble cocktails. The possible application for the acoustic heating of the inner solar atmosphere up to the transition region is shortly discussed together with the evaluation to confirm the theory by laboratory plasmas.
@@ I.2 A short discussion on the bulk viscosity
In the air we breath the energy transfer between translational and intramolecular degree of freedom is relatively slow and he first and second viscosity are comparable $\zeta/\eta\sim 1$. Perhaps Albert Einstein was the first who paid attention on the theory of bulk viscosity in partially dissociated gases [7]. The direct translation of this paper is Sound propagation in partially dissociated gases and in this sense, our study is a continuation of Einstein’s idea realized in partially ionized atomic gases with the main focus being the solar chromosphere.
@@
For the importance of the bulk viscosity $\zeta$ in the physics of plasmas we wish to point out the recent study by Istomin & Kustova [12] who studied transport coefficients and heat fluxes in non-equilibrium high-temperature flows with electronic excitation. They confirmed the intuitive conclusions by Einstein and calculated that chemical processes can lead to bulk viscosity domination $\zeta\gg\eta$. The authors of this study performed a non-equilibrium kinetic calculation of N and O ionized mixtures, while our study is focused on monoatomic plasmas close to local thermodynamic equilibrium (LTE). However, their conclusions and the results in our work are qualitatively the same. Their conclusion is that the bulk viscosity coefficients are strongly affected by the electronic excitation which is quantitatively similar to ours for the influence of the ionization-recombination processes to the bulk viscosity and degree of ionization $\alpha$. Istomin and Kustova [12, Figure 3] in their study demonstrate that the bulk viscosity has a broad maximum as a function of the temperature. Moreover bulk viscosity significantly exceeds the shear one analogously to the atomic plasma which is the subject in the present study. For the hydrogen-helium cocktail in the present article we have derived for the dispersionless bulk viscosity $\zeta_{0}\propto(1-\alpha)\alpha$, and at fixed mass density $\rho$, the ionization degree $\alpha(T)$ monotonously increases between 0 and 1, and again as a function of the temperature we have a broad maximum.
Istomin & Kustova [12] emphasize how important is for calculation of bulk viscosity is to have reliable data for the temperature dependence of the rates of the chemical reactions. The main our motivation is that for a hydrogen like atom since more than 70 years we have a reliable formula for near threshold ionization cross=section which gives the opportunity to use analytical formula for the rate of ionization $\beta$.
In Ref [12] dense molecular plasma is studied, which arises in the Earth’s atmosphere at re-entrance of spacecrafts, for example. While the results of our work find application in the partially ionized solar chromosphere and its sound wave heating. Although these studies focus on different physical scenarios, high vs low dense plasma, the common conclusion is however the same, up to now the bulk viscosity has not been in the focus of plasma physics, despite many interesting applications. This is also noted in Ref. [12, Sec. Introduction]: However, in the majority of studies (except Refs. 10 and 12–14) atomic species are considered as structure-less particles. Under this assumption, no internal heat conductivity and bulk viscosity coefficients appear for atomic species, both neutral and ionized, which can cause the loss of accuracy in the transport terms evaluation.
@@ II.2 The Saha equation
Often the corresponding concentration is used [15, Sec. 104] and it is straightforward to change the notations
@@
Within the framework of the up to now introduced notations, the Saha equation for the equilibrium concentrations reads in its standard form [15, Eq. (104.2)]
@@ II.3 Thermodynamic variables
Knowing these thermodynamic potentials, we can easily calculate heat capacity per unit mass from the derivatives [15, Sec. 16]
$\mathcal{C}_{p}$ $\equiv\left(\dfrac{\partial w}{\partial T}\right)_{\!\!p},$ (19)
@@
The sound velocity at evanescent frequency can be calculated by the standard thermodynamic derivatives [13, Eq. (A9)]
@@
In the formulae above, the overline denoting thermal equilibrium values have been omitted again. The velocity $c_{0}$ is the sound velocity at low frequency for which during the adiabatic compression of the plasma, the ionization degrees can follow the equilibrium Saha equation.
@@ III Wannier ridge resonances and Wannier near threshold ionization cross-section
It is remarkable that the exact solution for the scattering in the Coulomb potential coincides with both Born and classical cross-section approximations. The physics is however, absolutely different. At high energies we have slightly perturbed plane waves while the classical approximation is applicable when the velocity $v$ of the scattering electron is much smaller than the Bohr velocity
@@
The numerical values we use $\mathrm{w}\approx 1.18\sim 1$ and $C_{\mathrm{W}}\approx 2.7$ are obtained from the experimental study [20, Fig. 6]. We consider as a relevant discussion for plasma physics to explain that the Wannier formula is applicable within an acceptable accuracy for all $s$-electrons ionized by electron impact. The energy of the classical motion is just the quantum ionization potential.
Performing [21, Eq. (5)] the integration over equilibrium Maxwell distribution of electrons
@@ IV.4 Generalized polytropic index for cold plasmas
As test for the working programs we have to check that for very low and very high frequencies the polytropic index is real only. For high frequencies it reaches the atomic value $\gamma_{a}$ corresponding to negligible chemical reactions. For low frequencies we have a sophisticated thermodynamic expression Eq. (23) [13, Eq. (8)]
@@ V.1 Mandelstam-Leontovich and Cole-Cole approximation
see also [21, Eq. 99].
The excellent accuracy of the ML fit with fitting parameter time-constant $\tau$ gives a simple interpretation of the both low- and high-frequency limits; we have a simple Padé approximation. We can consider the oscillation of the density as a perturbation and the pressure as a linear response. According to the causality principle, the generalized susceptibility $\hat{p}/\hat{\rho}\propto\hat{\gamma}(\omega)$ must be analytical in the upper semi-plane of the complex frequency $\Im(\omega)>0$ and just along the imaginary semi-axis $\omega=\mathrm{i}\tilde{s}$ with $\tilde{s}>0$ the compressibility must be real and positive. In the lower imaginary semi-plane $\hat{\gamma}_{\mathrm{rel}}^{\mathrm{{ML}}}(\omega)$ has a zero at $\omega=-\mathrm{i}/\tau$ and a pole at $\omega=-\mathrm{i}\Gamma_{0,\infty}/\tau$ in agreement with the general causality principle [15, Sec. 123 Generalized susceptibility].
@@
$A=\exp\left(-\int_{0}^{h_{\mathrm{TR}}}2k^{\prime\prime}(x)\mathrm{d}x\right)$ (116)
@@
$\zeta_{0}=-\overline{p}\lim_{\omega\rightarrow 0}\frac{\gamma^{\prime\prime}(\omega)}{\omega}=-\overline{p}\left.\frac{\mathrm{d}\gamma^{\prime\prime}}{\mathrm{d}\omega}\right|_{\omega=0}.$ (125)
@@
VI Hydrogen-helium (alkali-noble) cocktail analytical solution for the thermodynamics and kinetics
@@
$\overline{\mathcal{N}}_{\mathrm{tot}}=1+\overline{\alpha}+\overline{a}_{\mathrm{He}},$ (144)
@@
And for the pressure in the approximation of ideal gas we have
$p=n_{\mathrm{tot}}T=\mathcal{N}_{\mathrm{tot}}n_{\rho}T=(1+\alpha+\overline{a}_{\mathrm{He}})n_{\rho}T.$ (146)
The Saha equation can be written as
$\alpha(\rho,T)=f(\nu)\equiv\alpha_{\mathrm{H}}\equiv\dfrac{2}{1+\sqrt{1+4\nu}},$ (147)
$\nu\equiv\dfrac{n_{\rho}}{n_{\mathrm{{}_{S}}}}=\frac{1-\alpha}{\alpha^{2}},\qquad\dfrac{1}{\sqrt{1+4\nu}}=\dfrac{\alpha}{2-\alpha}$ (148)
$n_{\mathrm{{}_{S}}}(T)\equiv n_{q}\mathrm{e}^{-\iota},\quad n_{q}(T)=\left(\dfrac{mT}{2\pi\hbar^{2}}\right)^{\!\!3/2},\quad\iota\equiv\dfrac{I}{T},$
@@
where $m$ is the electron mass and $n_{\rho}=\rho/M^{*}$, Its solution for the hydrogen degree of ionization $\overline{\alpha}(T,\overline{n}_{\rho})$ is given in Ref. [21]. Actually, this is an applicable approximation for the solar chromosphere but the derived formulas are applicable for any alkali-noble cocktail, Na-Ne plasma or arbitrary cold pure element plasma.
VI.1 Thermodynamics of the cold binary plasma
For the enthalpy $w$ and internal energy $\varepsilon$ per unit mass we have [25, 13, Eqs. (11-12), Eq. (6)]
$w(\rho,T)$ $=\dfrac{1}{M^{*}}\left(c_{p}\mathcal{N}_{\mathrm{tot}}T+I\alpha\right),\qquad c_{p}=\dfrac{5}{2},$ (149)
$\varepsilon(\rho,T)$ $=\dfrac{1}{M^{*}}\left(c_{v}\mathcal{N}_{\mathrm{tot}}T+I\alpha\right),\qquad c_{v}=\dfrac{3}{2}.$ (150)
Here we wish to emphasize that magnetic field has zero influence on the thermodynamics of the solar atmosphere which in a classical gas. This statement is known as Bohr–van Leeuwen [26] theorem. For the derivatives we have
$f^{\prime}(\nu)$ $=\mathrm{d}_{\nu}f(\nu)=-\dfrac{4}{\left(1+\sqrt{1+4\nu}\right)^{\!2}\sqrt{1+4\nu}}$
$=\dfrac{-f^{2}(\nu)}{\sqrt{1+4\nu}}=\dfrac{-f^{3}}{2-f}=-\dfrac{\alpha^{3}}{2-\alpha}=\dfrac{\mathrm{d}\alpha}{\mathrm{d}\nu},$ (151)
$\nu f^{\prime}(\nu)$ $=-\dfrac{(1-\alpha)\alpha}{2-\alpha}=-\mathcal{D},\quad\mathcal{D}\equiv\dfrac{(1-\alpha)\alpha}{2-\alpha},$ (152)
$\left(\dfrac{\partial\alpha}{\partial T}\right)_{\!\!\rho}$ $=-\dfrac{(c_{v}+\iota)}{T}\nu f^{\prime}(\nu)=\dfrac{(c_{v}+\iota)}{T}\dfrac{(1-\alpha)\alpha}{2-\alpha}$
$=\dfrac{(c_{v}+\iota)}{T}\mathcal{D},$ (153)
$\left(\dfrac{\partial\alpha}{\partial\rho}\right)_{\!T}$ $=\nu f^{\prime}(\nu)/(M^{*}n_{\mathrm{{}_{S}}})=-\dfrac{1}{M^{*}n_{\mathrm{{}_{S}}}}\dfrac{(1-\alpha)\alpha}{2-\alpha}$
$=-\dfrac{1}{M^{*}n_{\mathrm{{}_{S}}}}\mathcal{D}.$ (154)
In such a way we obtain for the heat capacity the explicit expressions
$\mathcal{C}_{v}$ $\equiv\left(\dfrac{\partial\varepsilon}{\partial T}\right)_{\!\!\rho}=\dfrac{1}{M^{*}}\left[c_{v}\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)^{2}\mathcal{D}\right]$ (155)
$=\dfrac{1}{M^{*}}\left[c_{v}(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)^{2}\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]$ (156)
$=T\left(\dfrac{\partial s}{\partial T}\right)_{\!\!\!\rho},$
where $s$ is the entropy per unit mass.
For derivatives of the pressure Eq. (146) we have
$\left(\dfrac{\partial p}{\partial T}\right)_{\!\!\rho}$ $=\left[\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)\mathcal{D}\right]n_{\rho}$ (157)
$=\left[(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]n_{\rho},$
$\left(\dfrac{\partial p}{\partial\rho}\right)_{\!\!T}$ $=\dfrac{T}{M^{*}}\left[\mathcal{N}_{\mathrm{tot}}-\mathcal{D}\right]$ (158)
$=\dfrac{T}{M^{*}}\left[(1+\alpha+\overline{a}_{\mathrm{He}})-\dfrac{(1-\alpha)\alpha}{2-\alpha}\right].$
Heat capacity per constant volume and unit mass $\mathcal{C}_{v}$ can be obtained by the general formula [15, Eq. (16.10)] in which volume is for unit mass $\mathcal{V}\equiv 1/\rho$
$\Delta\mathcal{C}$ $\equiv\mathcal{C}_{p}-\mathcal{C}_{v}=-T\dfrac{(\partial p/\partial T)_{\mathcal{V}}^{2}}{(\partial p/\partial\mathcal{V})_{T}}=\dfrac{T}{\rho^{2}}\dfrac{(\partial p/\partial T)_{\rho}^{2}}{(\partial p/\partial\rho)_{T}}$
$=\dfrac{\left[\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)\mathcal{D}\right]^{2}}{M^{*}\left[\mathcal{N}_{\mathrm{tot}}-\mathcal{D}\right]}$ (159)
$=\dfrac{\left[(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]^{2}}{M^{*}\left[(1+\alpha+\overline{a}_{\mathrm{He}})-\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]},$
and
$\mathcal{C}_{p}=$ $\dfrac{\left[c_{v}\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)^{2}\mathcal{D}\right]}{M^{*}}+\dfrac{\left[\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)\mathcal{D}\right]^{2}}{M^{*}\left[\mathcal{N}_{\mathrm{tot}}-\mathcal{D}\right]}$
$=$ $\dfrac{\left[c_{p}\mathcal{N}_{\mathrm{tot}}+(c_{v}c_{p}+2c_{p}\iota+\iota^{2})\mathcal{D}\right]\mathcal{N}_{\mathrm{tot}}}{(\mathcal{N}_{\mathrm{tot}}-\mathcal{D})M^{*}}$ (160)
$=$ $\dfrac{1}{M^{*}}\left[c_{v}(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)^{2}\frac{(1-\alpha)\alpha}{2-\alpha}\right]$
$+\dfrac{\left[(1+\alpha+\overline{a}_{\mathrm{He}})+(c_{v}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]^{2}}{M^{*}\left[(1+\alpha+\overline{a}_{\mathrm{He}})-\dfrac{(1-\alpha)\alpha}{2-\alpha}\right]}$ (161)
$=T\left(\dfrac{\partial s}{\partial T}\right)_{\!\!p}.$ (162)
These results generalize the previously obtained heat capacities for pure hydrogen [25].
Analogously for the derivatives of enthalpy
$\left(\dfrac{\partial w}{\partial T}\right)_{\!\!\rho}$ $=\dfrac{1}{M^{*}}\left[c_{p}\mathcal{N}_{\mathrm{tot}}+(c_{v}+\iota)(c_{p}+\iota)\mathcal{D}\right]$ (163)
$=\left[(1+\alpha+\overline{a}_{\mathrm{He}})c_{p}+(c_{v}+\iota)(c_{p}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right],$
$\left(\dfrac{\partial w}{\partial\rho}\right)_{\!\!T}$ $=-\dfrac{T}{M^{*}}(c_{p}+\iota)\dfrac{\mathcal{D}}{M^{*}n_{\rho}}$ (164)
$=-\dfrac{T}{M^{*}}\dfrac{1}{M^{*}n_{\rho}}\left[(c_{p}+\iota)\dfrac{(1-\alpha)\alpha}{2-\alpha}\right].$
Then the Jacobian [13, Eq. (9)] with dimension of reciprocal mass
$\mathcal{J}\equiv\frac{\partial(w,p)}{\partial(T,\rho)}=\left(\frac{\partial w}{\partial T}\right)_{\!\!\rho}\left(\frac{\partial p}{\partial\rho}\right)_{\!\!T}-\left(\frac{\partial w}{\partial\rho}\right)_{\!\!T}\left(\frac{\partial p}{\partial T}\right)_{\!\!\rho}$ (165)
after some algebra reads
$\mathcal{J}=\frac{T\mathcal{N}_{\mathrm{tot}}}{(M^{*})^{2}}\left\{c_{p}\mathcal{N}_{\mathrm{tot}}+\left[(c_{v}+\iota)(c_{p}+\iota)+\iota\right]\mathcal{D}\right\}=\frac{(1\!+\!\alpha\!+\!\overline{a}_{\mathrm{He}})T}{(M^{*})^{2}}\left[c_{p}(1\!+\!\alpha\!+\!\overline{a}_{\mathrm{He}})+(c_{v}c_{p}+2c_{p}\iota+\iota^{2})\frac{(1-\alpha)\alpha}{2-\alpha}\right].$ (166)
In such a way we obtained that for the 2-component cocktail
$\mathcal{J}=\frac{T}{M^{*}}(\mathcal{N}_{\mathrm{tot}}-\mathcal{D})\mathcal{C}_{p}.$ (167)
Representing the sound velocity at evanescent frequency when ionization degree follows it equilibrium values determined by Saha equation we have
$c_{0}^{2}\equiv\left(\frac{\partial p}{\partial\rho}\right)_{\!\!s}=\gamma_{0}\frac{p}{\rho},$ (168)
The general result [13, Eq. (9)]
$\gamma_{0}=\frac{\rho}{p}\cdot\frac{\mathcal{J}}{\mathcal{C}_{p}}$ (169)
after substitution $\mathcal{J}$ from Eq. (167) gives
$\gamma_{0}$ $=$ $\left(1-\dfrac{\mathcal{D}}{\mathcal{N}_{\mathrm{tot}}}\right)\dfrac{\mathcal{C}_{p}}{\mathcal{C}_{v}}$
$=$ $\left[1-\dfrac{(1-\alpha)\alpha}{(2-\alpha)(1+\alpha+\overline{a}_{\mathrm{He}})}\right]\dfrac{\mathcal{C}_{p}}{\mathcal{C}_{v}}.$
This illustrates how the ionization recombination processes break the general relations derived for constant chemical compounds. For solar plasma it is not so strongly expressed but for interstellar plasma where $\iota=I/T\gg 1$ it is possible $\iota^{2}\mathcal{D}$ to dominate, hence
$M^{*}\mathcal{C}_{v}$ $\approx$ $\iota^{2}\mathcal{D}\gg 1,$ (171)
$M^{*}\mathcal{C}_{p}$ $\approx$ $\dfrac{\iota^{2}\mathcal{D}}{1-\mathcal{D}/\mathcal{N}_{\mathrm{tot}}}\gg 1,$ (172)
confer Refs. [25, 21, Fig. 3 and Fig. 5]. In the opposite case of negligible influence of the ionization processes, $\iota\ll 1$ we have the test for the programming of the heat capacities
$M^{*}\mathcal{C}_{v}\approx c_{v}\mathcal{N}_{\mathrm{tot}},\;M^{*}\mathcal{C}_{p}\approx c_{p}\mathcal{N}_{\mathrm{tot}},\;M^{*}(\mathcal{C}_{p}-\mathcal{C}_{v})=\mathcal{N}_{\mathrm{tot}}.$ (173)
The general formula for pure hydrogen plasma was solved in the early stages of the present study [21, Eq. (96)] and [25, Eq. (17)]
where $\iota\equiv I/T$, $I=13.6$ eV is the hydrogen ionization potential, $\alpha\equiv\overline{\alpha}$ is the equilibrium ionization degree
$\alpha\equiv\frac{n_{e}}{n_{p}+n_{0}}=\frac{n_{e}}{n_{\rho}}\in(0,1)\,$, (175)
$n_{e}=n_{p}=\alpha n_{\rho},\quad n_{0}=(1-\alpha)\,n_{\rho},\quad$ (176)
$n_{\mathrm{tot}}=n_{e}+n_{p}+n_{0}=(1+\alpha)\,n_{\rho},$ (177)
where $n_{e}$, $n_{p}$ and $n_{0}$ are the concentrations of electrons, protons and neutral atoms respectively, and we denote the sum of the concentrations of the protons and the neutrals by $n_{\rho}=n_{0}+n_{p}$. This formula can be generalized for arbitrary low temperature cocktail of alkaline metal and noble gas [13, Eq. (B1)], (for example Na-Ne plasma)
$\gamma_{0}=\frac{c_{p}[2+(2-\alpha)\overline{a}_{\mathrm{nbl}}]+(1-\alpha)\alpha(c_{p}+\iota)^{2}}{c_{v}[2+(2\!-\!\alpha)\overline{a}_{\mathrm{nbl}}]+(1\!-\!\alpha)\alpha\left[(c_{v}+\iota)^{2}+c_{v}\right]},$ (178)
where
VI.2 Kinetics of the cold binary plasma